You are analyzing the stock of Pizza Hub, an online pizza delivery company. Your estimate of the cost of equity is 16 percent. The most recent dividend is $0.20 per share. You expect dividends to grow 15 percent per year for three years and 8 percent per year thereafter. What is your estimate of the stock price?
Answer #1

Dividend in year 1, D1= D0 x (1+g) = $ 0.2 x (1+0.15) = $ 0.2 x 1.15 = $ 0.23

Dividend in year 2, D2= D1 x (1+g) = $ 0.23 x (1+0.15) = $ 0.23 x 1.15 = $ 0.2645

Dividend in year 3, D3= D2 x (1+g) = $ 0.2645 x (1+0.15) = $ 0.2645 x 1.15 = $ 0.304175

Dividend in year 4, D4= D3 x (1+g) = $ 0.304175 x (1+0.08) = $ 0.304175 x 1.08 = $ 0.328509

As per Dividend Discount Model stock price in year 3 can be computed as:

P3 = D4/(r – g)

r = cost of capital = 16 % or 0.16

g = Constant growth rate of dividend = 8 % or 0.08

     P 3 = $ 0.328509/ (0.16 – 0.08)

     = $ 0.328509/0.08

      = $ 4.1063625

Current stock price P0 = D1 / (1+r) + D2 / (1+r) 2 + D3 / (1+r) 3 + P3 / (1+r) 3

  = $ 0.23 / (1+0.16) + $ 0.2645 / (1+0.16) 2 + D3 / (1+0.16) 3 + P3 / (1+0.16) 3

= $ 0.23 / (1.16) + $ 0.2645 / (1.16) 2 + $ 0.304175/ (1.16) 3 + $ 4.1063625/ (1.16) 3

= $ 0.23 /1.16 + $ 0.2645 / 1.3456 + $ 0.2645/ 1.560896 + $ 4.1063625/ 1.560896

= $ 0.198275862068966 + $ 0.196566587395957 + 0.19487204784944 + 2.63077264596744

= $ 3.22048714328181 or $ 3.22

Current stock price of Pizza Hub is $ 3.22